OpenStudy (ashley1nonly):

check my work Determine the rate law and the value of k for the following reaction using the data provided. 2 N2O5(g) → 4 NO2(g) + O2(g) [ N2O5]i (M) Initial Rate (M-1s-1) 0.093 4.84 x 10-4 0.084 4.37 x 10-4 0.224 1.16 x 10-3 A) Rate = 5.6 × 10-2 M-1s-1[N2O5]2 B) Rate = 6.0 × 10-1 M-2s-1[N2O5]3 C) Rate = 1.6 × 10-3 M1/2s-1[N2O5]1/2 D) Rate = 1.7 × 10-2 M-1/2s-1[N2O5]3/2 E) Rate = 5.2 × 10⁻3 s-1[N2O5] Answer: E

OpenStudy (ashley1nonly):

the rate equal rate=k[a] rate=k[N2O5]^2 (4.84*10^-4)=k(0.093) divide by .093 5.6*10^-2 = k the units for k m/s=k(m)^2 m/s * 1/m^2 k= m^-1s^-1 k= 5.6*10^-2 m^-1s^-1 what did i do wrong to not get E

OpenStudy (ashley1nonly):

i think it is just an error

OpenStudy (ashley1nonly):

@Preetha

OpenStudy (ashley1nonly):

@mayankdevnani

OpenStudy (ashley1nonly):

@abb0t